Live GMAT classes with Target Test Prep start Nov 21 and Nov 22. First class free.

See classes
Target Test Prep · GMAT

Choose how you want to prepare

Learn live with an expert or move at your own pace. Every option includes the complete TTP study system.

Newsweek Readers’ Choice Awards 2026 winner: Target Test Prep, #1 Test Prep

Target Test Prep was voted #1 Best Test Prep in the 2026 Newsweek Readers’ Choice™ Awards.

★★★★★5.0563 reviews
GMATLiveTeachStarts Nov 22
Josh Braslow, Target Test Prep GMAT expert
LIVE ONLINE CLASSES

Get Ready for GMAT Test Day Faster with Live Online Classes

with Josh Braslow, 20+ years in test prep

Nov 22 · Josh Braslow
Sun · 9:30 AM to 12:30 PM ET
Included
40 hours of live online classes + 6 months of TTP OnDemand
  • Attend the first class for free
  • Every class is recorded, so you never fall behind
View classes & enroll
Limited seats availableTarget Test Prep
EALiveTeachOnDemandStart anytime
EXECUTIVE ASSESSMENT

Target Test Prep EA OnDemand

Self-paced EA prep. Study on your schedule.

EXECUTIVE ASSESSMENT

with

165+ EA score guarantee
$05-day trial no automatic billing
Schedule
Included
40 hours of live online classes plus six months of access to the complete TTP EA OnDemand course.
  • 165+ EA Score Guarantee
  • 4,100+ Quant, Verbal, and Integrated Reasoning practice questions
  • 400+ hours of in-depth video lessons
  • 3,000+ step-by-step video solutions
View EA class & enroll Start free 5-day trial
Limited cohort · enrollment openTrial includes full course accessTarget Test Prep
GMATOnDemandStart anytime
SELF-PACED MASTERCLASS

Target Test Prep GMAT OnDemand

Complete access from day one. Study on your schedule.

715+ score guarantee
$0to start then $127/mo
  • Personalized study plan and analytics
  • Thousands of lessons and practice questions

probability

Expert replies
by lakshnachadha » Sun Dec 05, 2010 6:53 am
A bag contains 3 white balls, 3 black balls & 2 red balls. One by one three balls are drawn out without replacement. What is the probability that the third ball is red?
Join the discussion
Source: — Problem Solving |

by shovan85 » Sun Dec 05, 2010 9:31 am
lakshnachadha wrote:A bag contains 3 white balls, 3 black balls & 2 red balls. One by one three balls are drawn out without replacement. What is the probability that the third ball is red?
3W 3B 2R

In order to get a Red in third pick, the two red balls should not have been picked in first two picks.

[P(Getting Red in first pick) AND P (Getting Something other color in second pick)] OR [P(Getting Something other color in first pick) AND P (Getting Red in second pick)]

= 2/8 * 6/7 + 6/8 * 2/7

= 3/7

Now, P(Getting Red in third pick) = 3/7 * 1/6 = 1/14

[P(Getting Something other color in second pick) AND P (Getting Something other color in second pick)]

= 6/8 * 5/7 = 15/28

Now, P(Getting Red in third pick) = 15/28 * 2/6 = 5/28.

Answer will be 1/14 + 5/28 = 7/28 = 1/4

Is it correct? Let me know.
If the problem is Easy Respect it, if the problem is tough Attack it
Join the discussion

by beat_gmat_09 » Sun Dec 05, 2010 9:34 am
Probability of drawing red ball on third attempt is same as probability of drawing red ball on first attempt.
Possible outcomes= 2
Total outcomes = 8
Probability = 2/8 = 1/4
Hope is the dream of a man awake
Join the discussion

by shovan85 » Sun Dec 05, 2010 9:44 am
beat_gmat_09 wrote:Probability of drawing red ball on third attempt is same as probability of drawing red ball on first attempt.
Possible outcomes= 2
Total outcomes = 8
Probability = 2/8 = 1/4
Phew!! Its that simple... hahaha ;)

Need rest badly. Good Night fellas.
If the problem is Easy Respect it, if the problem is tough Attack it
Join the discussion

by beat_gmat_09 » Sun Dec 05, 2010 9:46 am
shovan85 wrote:
beat_gmat_09 wrote:Probability of drawing red ball on third attempt is same as probability of drawing red ball on first attempt.
Possible outcomes= 2
Total outcomes = 8
Probability = 2/8 = 1/4
Phew!! Its that simple... hahaha ;)

Need rest badly. Good Night fellas.
C ya 2morrow.. good night..
Hope is the dream of a man awake
Join the discussion